OpenStudy (anonymous):

I have 7 8th grade math problems!

OpenStudy (anonymous):

I will give metals!

OpenStudy (anonymous):

How can i help?! =)

OpenStudy (anonymous):

okay I will post them!

OpenStudy (anonymous):

1.Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form. 2.Consider the following pair of equations: –x – y = –5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect? (3, 2) (3, –2) (–2, 3) (2, 3) 3.Consider the following set of equations: Equation R: –3y = –3x - 9 Equation S: y = x + 3 Which of the following best describes the solution to the given set of equations? No solution Infinite solutions Two solutions One solution 4.Consider the following system of equations: y = 5x + 6 y = –x – 7 Which description best describes the solution to the system of equations? Lines y = 5x + 6 and y = –x – 7 intersect the x-axis. Lines y = 5x + 6 and y = –x – 7 intersect the y-axis. Line y = 5x + 6 intersects line y = –x – 7. Line y = 5x + 6 intersects the origin. 5.Two lines, A and B, are represented by the following equations: Line A: 2x + y = 6 Line B: x + y = 4 Which statement is true about the solution to the set of equations? There is no solution. It is (4, 0). It is (2, 2). There are infinitely many solutions. 6.Consider the following set of equations: Equation A: y = –x + 5 Equation B: y = 6x – 2 Which of the following is a step that can be used to find the solution to the set of equations? –x + 5 = 6x - 2 –x = 6x + 2 –x + 5 = 5x –x – 2 = 6x + 5